Find Tge measure of <Q if m<R=7x-26 and m<P=5x-2
1
Expert's answer
2020-01-07T08:30:40-0500

Given

"\angle R =7x-26 \space and \space \angle P=5x - 2"


We know sum of the angles in a triangle is equal to "180^o"


So,

"\angle P + \angle Q + \angle R = 180"


"5x-2+ 7x-26+ \angle Q=180"

Now solve for "\angle Q"


"\angle Q=180-12x +28"

"\angle Q=208-12x"

Answer:


"Measure \space of \space Angle \space Q= \angle Q=(208-12x)^{\circ}"